OpenStudy (anonymous):

find the derivative using the product rule and the chain rule: (x²+1)(x²+2)^¼

OpenStudy (anonymous):

uv'+vu'

OpenStudy (lgbasallote):

where do you get confused?

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\]

OpenStudy (anonymous):

finding the derivative of (x²+1)

OpenStudy (anonymous):

where u is your inner function

OpenStudy (anonymous):

\[\frac{d}{dx}[x^n]=nx^{n-1}\]

OpenStudy (anonymous):

i know.. so it's just 1?

OpenStudy (anonymous):

x^n where your n is 2 in this case

OpenStudy (anonymous):

\[\frac{d}{dx}[k]=0\] where k is any constant

OpenStudy (anonymous):

I mean when you do the vu' part ... is it just (x²+2)^¼ (1)

OpenStudy (anonymous):

no since the derivative of x^2 = 2x

OpenStudy (anonymous):

because using the chain rule.. (x²+1) = 1 ... but the inside is 2x

OpenStudy (anonymous):

no need to d the chain rule for that

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

the chain function is when you have an outer and inner function

OpenStudy (anonymous):

OH! Ok thank you so much!

OpenStudy (anonymous):

such as \[ln(x^2+2)\] or \[(x^2+2)^2\]

OpenStudy (anonymous):

Yes I understand. Thank you so much.

OpenStudy (anonymous):

so whe you figure the answer out , post it

OpenStudy (anonymous):